# ICAT: Evaluating Completeness of Factual Information in Long-form Text Generation

### v1

@@ 7 Conclusions and Future Work
We presented ICAT, a comprehensive framework for evaluating topic coverage in LLM-generated text. Through extensive experimentation using the ClueWeb09 dataset, we demonstrated the framework’s effectiveness across different evaluation scenarios, with our best method achieving strong correlation with human judgments. The modular architecture of ICAT enables flexible adaptation to various evaluation requirements, from manual to automatic approaches for aspect identifications and alignment. Our results highlighted several key findings: (1) the importance of sophisticated coverage models in improving evaluation accuracy, (2) the viability of automatic evaluation approaches that maintain comparable performance to methods requiring ground truth annotations, and (3) the framework’s ability to provide meaningful assessments across different LLM architectures and scales.
In future work, individual components of our ICAT could be improved in terms of effectiveness and efficiency. Additionally, exploring the relationship between model size, evaluation accuracy, and computational efficiency could provide valuable insights for practical applications. Last but not least, the potential bias introduced by using the same or similar LLM when generating ground-truth aspects should be investigated. By using our metric, other works can explore methods for optimizing LLMs to produce more comprehensive outputs.
8 Limitations
@@
Third, there is a potential source of bias when using the same or similar LLM architecture both for generating query aspects and for producing responses for evaluation. This circular dependency might lead to artificially inflated performance metrics if the evaluated model shares similar biases or knowledge patterns with the model used for aspect generation. Future research should investigate the extent of this potential bias and explore methods to mitigate it, such as using diverse model architectures or more comprehensive human-curated aspects for evaluation.
9 Acknowledgments
@@ A.1 Atomic Claim Generation
Figure 2 illustrates our atomic claim generation process with a concrete example, showing how a complex paragraph is broken down into individual atomic claims while preserving key contextual information.
@@ A.2 Model Performance Analysis
Table 4 presents detailed evaluation results using Llama-3.1-8B as the coverage model. The results demonstrate consistent performance patterns across different retrieval methods (corpus-based vs web search) and highlight the trade-offs between factuality ($S_{F}$) and coverage ($S_{C}$) scores for various models.
@@
Web Search 0.376 0.570 0.411 0.570 0.411
Table 4: Evaluation results for ICAT-S and ICAT-A (Llama-3.1-8B Coverage Model)
Table 5 analyzes the impact of different retrieval models on the correlation between ICAT variants and human judgments. The results show that more sophisticated dense retrieval models (Snowflake-Arctic-Embed-L) consistently outperform traditional BM25 across all ICAT variants, with improvements particularly notable in ICAT 2 and 3.
@@ A.3 Prompting Details
Figures 4, 5, and 6 provide the detailed prompts used for subtopic coverage assessment, atomic claim generation, and topic generation respectively. These carefully designed prompts are crucial for ensuring consistent and high-quality outputs from the language models used in our framework.
@@ A.4 Human Evaluation Interface
Figure 7 shows our human annotation interface used in the evaluation study. The interface allows annotators to highlight supporting evidence for different aspects of the query in the LLM response, enabling fine-grained assessment of topic coverage. The visualization of multiple annotators’ highlights helps demonstrate the level of agreement in human judgments.
@@
ICAT-A BM25 0.483 0.449 0.374
Snowflake-Arctic-Embed-M 0.500 0.459 0.382
Snowflake-Arctic-Embed-L 0.512 0.468 0.389
Table 5: Impact of different retrieval models on correlation with Human Judgments using Llama-3.1-70B for topic-claim alignment.
Figure 4: Subtopic coverage prompt
Figure 5: Claim generation prompt
Figure 6: Topic generation prompt

### v4

@@ 7 Conclusions and Future Work
We presented ICAT, a comprehensive framework for evaluating topic coverage in LLM-generated text. Through extensive experimentation using the ClueWeb09 dataset, we demonstrated the framework’s effectiveness across different evaluation scenarios, with our best method achieving strong correlation with human judgments. The modular architecture of ICAT enables flexible adaptation to various evaluation requirements, from manual to automatic approaches for aspect identifications and alignment. Our results highlighted several key findings: (1) the importance of sophisticated coverage models in improving evaluation accuracy, (2) the viability of automatic evaluation approaches that maintain comparable performance to methods requiring ground truth annotations, and (3) the framework’s ability to provide meaningful assessments across different LLM architectures and scales. In future work, individual components of our ICAT could be improved in terms of effectiveness and efficiency. Additionally, exploring the relationship between model size, evaluation accuracy, and computational efficiency could provide valuable insights for practical applications. Last but not least, the potential bias introduced by using the same or similar LLM when generating ground-truth aspects should be investigated. By using our metric, other works can explore methods for optimizing LLMs to produce more comprehensive outputs.
Acknowledgments
Limitations
@@ Acknowledgments
In addition, there is a potential source of bias when using the same or similar LLM architecture both for generating query aspects and for producing responses for evaluation. This circular dependency might lead to artificially inflated performance metrics if the evaluated model shares similar biases or knowledge patterns with the model used for aspect generation. Future research should investigate the extent of this potential bias and explore methods to mitigate it, such as using diverse model architectures or more comprehensive human-curated aspects for evaluation.
Last but not least, our current evaluation is limited to English-language content using web-based corpora. This narrow focus excludes evaluation of multilingual capabilities and limits the framework’s applicability to other languages and cultures. Additionally, the reliance on web corpora may not be suitable for domains requiring specialized knowledge bases or authoritative sources. Future work should explore extending ICAT to support multilingual evaluation and integration with domain-specific knowledge bases to broaden its applicability across different languages, cultures, and specialized fields.
@@ A.1 Atomic Claim Generation
Our atomic claim generation process transforms complex text into atomic, verifiable statements while preserving essential context, as shown in Figure 2. The process focuses on creating decontextualized, self-contained claims that each express a single verifiable fact. When breaking down complex sentences, the process maintains important qualifiers, conditions, and temporal information while replacing contextual references with their explicit referents. For instance, a complex sentence about coffee’s effects would be decomposed into separate claims about its components and their individual effects, with each claim being fully self-contained and independently verifiable.
@@ A.2 Model Performance Analysis
The performance analysis presented in Tables 5 and 10 reveals several important patterns in model behavior. When comparing the 8B and 70B versions of Llama-3.1 as coverage models, we observe that the 8B model generally produces higher raw coverage scores, while the 70B model demonstrates stronger correlation with human judgments. This suggests a trade-off between computational efficiency and evaluation accuracy. The retrieval model comparison shows consistent advantages for dense retrieval methods over traditional BM25, with the Snowflake-Arctic-Embed models showing particular strength in handling queries where simple lexical matching is insufficient. Web-based retrieval consistently produces higher factuality scores compared to corpus-based approaches across all tested models. Table 10 analyzes the impact of different retrieval models on the correlation between ICAT variants and human judgments. The results show that larger dense retrieval models (Snowflake-Arctic-Embed-L) consistently outperform traditional BM25 across all ICAT variants, with improvements particularly notable in ICAT 2 and 3.
@@
Web Search 0.752 0.570 0.648 0.570 0.648
Table 5: Comprehensive evaluation results comparing ICAT-S and ICAT-A variants using Llama-3.1-8B as the coverage model. Results show performance across different LLMs using both corpus-based and web-based retrieval methods, demonstrating consistent improvements in factuality and coverage scores when using web search compared to corpus-based retrieval. The table highlights how different retrieval methods affect both factuality ($S_{\text{fact}}$) and coverage scores across various model architectures.
@@ A.3 Prompting Details
The framework employs three specialized prompts, each carefully designed for its specific task. The subtopic coverage assessment prompt (Figure 5) requests the identification of covered subtopics from a given list, requiring evidence in the form of fact numbers that explicitly appear in the text. The prompt specifies a structured JSON output format where each line contains a topic ID and its supporting evidence. The atomic claim generation prompt (Figure 6) focuses on extracting decontextualized, self-explanatory fact sentences from the input text, emphasizing the importance of resolved pronouns and independent context. The topic generation prompt (Figure 7) elicits possible subtopics or related queries for a given query, requiring them to be ordered by importance and formatted as JSON objects, with a maximum of 10 topics.
@@ A.4 Human Evaluation Interface
The human evaluation interface shown in Figure 4 was developed to facilitate consistent and accurate topic coverage assessment. The interface presents the query and response in a split-screen layout, enabling annotators to highlight evidence for different topics using a color-coded system. To ensure annotation quality, we implemented minimum time requirements per response and included attention check questions. Annotators were guided to read responses completely before beginning their annotations and to identify text spans that provide direct evidence for topic coverage.
A.5 Component-level Evaluation
Atomic Claim Generation Quality. We evaluated the quality of atomic claim generation by comparing our fine-tuned Llama-3.1-8B model against the zero-shot Llama-3.1-70B model on 50 randomly sampled queries. Human annotators created ground-truth atomic facts for each query, and we assessed the generated claims using multiple metrics. To enforce local comparisons between atomic facts, we used a sliding window approach (window size ±3) when computing similarity metrics.
We evaluated using several metrics: (1) BERTScore333using microsoft/deberta-v3-base metrics for token-level similarity; (2) Semantic Precision/Recall/F1 computed using cosine similarities of sentence embeddings444using Alibaba-NLP/gte-base-en-v1.5, where precision measures how many generated claims have semantically similar ground-truth claims, recall measures coverage of ground-truth claims, and F1 is their harmonic mean; (3) Completeness, similar to semantic recall but with similarities binarized at threshold 0.8 to measure hard coverage; (4) traditional metrics (ROUGE, BLEU); and (5) Fact Precision, which measures the proportion of atomic claims that are properly formulated—specifically, claims that are decontextualized, self-contained, and verifiable in isolation without pronouns or missing context based on human judgements.
As shown in Table 6, the 70B model achieves higher scores on most semantic similarity metrics (e.g., BERTScore F1: 0.571 vs 0.499), suggesting better overall alignment with ground-truth facts. However, the fine-tuned 8B model demonstrates superior fact precision (0.838 vs 0.753), indicating better decontextualization and self-containment of claims—crucial properties for accurate retrieval-based grounding. This trade-off between semantic similarity and proper claim formulation justifies our choice of using a fine-tuned smaller model for efficient claim generation.
Metric Llama-3.1-8B Llama-3.1-70B Diff
BERTScore Precision 0.497 0.584 +0.088
BERTScore Recall 0.514 0.569 +0.055
BERTScore F1 0.499 0.571 +0.072
Semantic Precision 0.618 0.726 +0.109
Semantic Recall 0.808 0.825 +0.017
Semantic F1 0.681 0.761 +0.080
Completeness 0.537 0.593 +0.056
ROUGE-1 F 0.703 0.706 +0.003
ROUGE-2 F 0.555 0.563 +0.008
ROUGE-L F 0.599 0.610 +0.011
BLEU 0.379 0.391 +0.012
Fact Precision 0.838 0.753 -0.085
Table 6: Evaluation of atomic claim generation quality comparing fine-tuned 8B and zero-shot 70B models. The 8B model shows superior fact precision despite lower semantic similarity scores.
Topic Generation Quality. We assessed the relevance of automatically generated subtopics for the same 50 queries. Human annotators evaluated each generated subtopic as relevant or irrelevant to the query. Table 7 shows that both 8B and 70B models achieve remarkably high subtopic precision (0.966 and 0.958, respectively), demonstrating that even smaller models can effectively generate relevant query aspects when properly prompted. This validates our approach of using LLM-generated topics in ICAT-A when ground-truth aspects are unavailable.
Model Subtopic Precision
Llama-3.1-8B 0.966
Llama-3.1-70B 0.958
Table 7: Human evaluation of topic generation quality. Both models achieve high precision in generating relevant subtopics.
Coverage Assessment Accuracy. To evaluate the accuracy of our aspect-claim alignment module, we compared the coverage predictions of ICAT-S (using Llama-3.1-70B) against human judgments on the set of ground-truth topics. Table 8 shows strong agreement metrics, with high precision indicating that the topics identified as covered by our system are indeed covered according to human judgment, and good recall showing that most human-identified covered topics are captured by our system.
Coverage Metric Value
Coverage Precision 0.903
Coverage Recall 0.798
Coverage F1 0.835
Table 8: Accuracy of ICAT-S topic coverage predictions compared to human judgments using Llama-3.1-70B as the coverage model.
NLI Model Performance. Our claim grounding relies on a DeBERTa-based NLI model for efficiency. Table 9 shows the model’s performance across multiple NLI benchmarks. The strong performance on MNLI (0.903) and FEVER-NLI (0.777) demonstrates effectiveness on standard entailment and fact verification tasks. While performance on adversarial datasets (ANLI) is lower, as expected, the model still maintains reasonable accuracy even on the challenging ANLI-R3 subset (0.495). These results justify using this efficient model over more expensive LLM-based approaches for claim verification.
Dataset Accuracy
MNLI 0.903
FEVER-NLI 0.777
ANLI-all 0.579
ANLI-R3 0.495
Table 9: Performance of the DeBERTa-based NLI model used for claim grounding across multiple benchmarks.
@@ A.5 Component-level Evaluation
ICAT-A BM25 0.242 0.225 0.187
Snowflake-Arctic-Embed-M 0.251 0.236 0.191
Snowflake-Arctic-Embed-L 0.256 0.234 0.195
Table 10: Comparative analysis of retrieval model impact on ICAT variants’ correlation with human judgments, using Llama-3.1-70B for topic-claim alignment. Results demonstrate the superiority of dense retrieval models (Snowflake-Arctic-Embed) over traditional BM25 across all ICAT variants, with the largest improvements seen in ICAT-S and ICAT-A. The analysis includes three correlation metrics (Pearson’s, Spearman’s, and Kendall’s) to provide a comprehensive view of alignment with human assessments across different statistical measures.
Figure 5: Subtopic coverage prompt used to identify which subtopics are covered in a given text. The prompt instructs the model to analyze a list of atomic claims and return covered subtopics with supporting evidence in a structured JSON format. Each response must include topic IDs and corresponding fact numbers that explicitly appear in the text.
Figure 6: Claim generation prompt designed to extract atomic factual statements from input text. The prompt emphasizes the importance of creating decontextualized, self-explanatory sentences with resolved pronouns and independent context, ensuring each generated claim can stand alone as a verifiable statement.
Figure 7: Topic generation prompt used to automatically generate diverse subtopics for a given query. The prompt requests possible subtopics or related queries ordered by importance, with output formatted as JSON objects and limited to a maximum of 10 topics. This automated approach enables reference-free evaluation when ground-truth topics are unavailable.